If marginal propensity to save is 0.6 and the government decreases its spending by R50 billion, then south africa ' s real GDP or total income will
1
Expert's answer
2019-04-15T10:28:22-0400

MPC=1-MPS=1-0.6=0.4

m=1/MPC=1/0.4=2.5

Total income will decreases:

m*R

2.5*50=125 billion